What is the sum of a 30–term arithmetic sequence where the first term is 73 and the last term is –101?
Answer by jim_thompson5910(35256)


The formula is Sn = (n*(a1 + an))/2

In this case,

a1 = 73

an = -101

n = 30

Plug all this into the formula above to get

Sn = (n*(a1 + an))/2

Sn = (30*(73 + -101))/2

Sn = (30*(-28))/2

Sn = (-840)/2

Sn = -420

So the answer is -420